Please help me how to solve An alcoholic drink claims 12% alcohol b volume. Calculate the volume in ml of alcohol present in a 0.8L wine
Found 2 solutions by math_tutor2020, josgarithmetic:
Answer by math_tutor2020(3817)


Answer: 0.096 liters

Work Shown:
12% of 0.8 liters = 0.12*0.8 = 0.096 liters

Note: 0.096 liters = 0.096*1000 = 96 mL

Answer by josgarithmetic(39618)

----------------------------------------------------------
An alcoholic drink claims 12% alcohol b volume. Calculate the volume in ml of alcohol present in a 0.8L wine
---------------------------------------------------------

Two things not necessarily the same thing.
First sentence is about "alcoholic drink". Second sentence is about "wine". Bad whole problem description.